# Lambda Expected Shortfall

### v2

@@ 1 Introduction
Beyond the foundational definition and properties of $\Lambda$-ES, which are the topics of Section 3, we proceed to conduct a comprehensive analysis of this new class of risk measures. In Section 4, we obtain a dual representation of $\Lambda$-ES (Theorem 3), offering deeper insights into its theoretical structure and connections to quasi-convex cash-subadditive risk measures. In Section 5, we explore the properties of $\Lambda$-ES in optimization problems, both as an objective function to minimize and as a constraint to impose, and analyze various forms of convexity in relevant reformulations of ES optimization problems. In Section 6, results are naturally extended to the space $L^{1}$ of integrable random variables, sometimes under slightly stronger assumptions. Section 7 concludes the paper. Some alternative potential formulations for $\Lambda$-ES are discussed in Appendix A, demonstrating why our proposed definition is the most robust, theoretically consistent, and desirable for risk management applications.
@@ 2.1 Risk measures
We start with risk measures that are used to quantify risks. A risk measure is a mapping $\rho:\mathcal{X}\to\overline{\mathbb{R}}$, where $\mathcal{X}$ is the space of random variables where $\rho$ is defined. Below, we list several common properties that a risk measure may satisfy, and their financial interpretation is well documented in the literature (e.g., Artzner et al., 1999; Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002). The risk measure $\rho$ is called a monetary risk measure if it satisfies
@@
  • $L^{1}$-continuity: $\rho(X_{n})\to\rho(X)$ for all $X,X_{1},X_{2}\dots\in\mathcal{X}$ and $X_{n}\stackrel{{\scriptstyle L^{1}}}{{\to}}X$ as $n\to\infty$.
@@ 5.2 Optimization with Lambda ES
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. We demonstrate general ideas of solving optimization problems with $\Lambda$-ES as a constraint or an objective. For $n\in\mathbb{N}$, let $\mathbf{L}=(L_{1},\dots,L_{n})\in(L^{\infty})^{n}$ represent a vector of losses, $\bm{\theta}\in\Theta$ represent a decision variable, where $\Theta$ is a convex set of actions, and $f:\Theta\times\mathbb{R}^{n}\to\mathbb{R}$ represent a loss function. For a typical example in finance, we can use $\mathbf{L}$ as the vector of losses from multiple assets and $\bm{\theta}$ as a portfolio weight vector.
First, we are interested in a problem where the decision maker minimizes an objective risk measure $\rho:L^{\infty}\to\overline{\mathbb{R}}$ of the aggregate loss $f(\bm{\theta},\mathbf{L})$, guaranteeing that the $\Lambda$-ES of the total loss does not exceed a pre-specified value $\ell\in\mathbb{R}$. Namely, we consider the following optimization problem:
@@
The following result provides a possible direction to simplify the problem above.
@@
Proposition 4 implies that an optimization problem with a $\Lambda$-ES constraint of level $\ell\in\mathbb{R}$ can be equivalently converted to a problem with the constraint on $\mathrm{ES}_{\Lambda(\ell)}$ of the same level $\ell$. ES-constrained optimization problem has been studed extensively in the literature (see e.g., Krokhmal et al., 2002).
Another natural question to study is how to minimize $\Lambda$-ES as an objective:
@@
The result below provides general insights into solving $\Lambda$-ES-based optimization problems. In specific problems, we may also consider some constraints along with the problem (25), which does not change the nature of our equivalence result. Define the mapping $\widetilde{T}_{\Lambda}:\Theta\times\mathbb{R}\times\mathbb{R}\times(L^{\infty})^{n}\to\overline{\mathbb{R}}$ as
$\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})=\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(f(\bm{\theta},\mathbf{L})-a)_{+}\right]\vee x.$ (26)
@@
For a right-continuous decreasing function $\Lambda:\mathbb{R}\to[0,1]$,
$\min_{\bm{\theta}\in\Theta}\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L}))=\min_{(\bm{\theta},a,x)\in\Theta\times\mathbb{R}\times\mathbb{R}}\left\{\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(f(\bm{\theta},\mathbf{L})-a)_{+}\right]\vee x\right\}.$ (27)
Moreover, for $\widetilde{T}_{\Lambda}$ defined in (26),
@@
    $\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is convex in $a\in\mathbb{R}$ for all $(\bm{\theta},x,\mathbf{L})\in\Theta\times\mathbb{R}\times(L^{\infty})^{n}$.
@@
    if in addition, $\Lambda$ is not constantly $1$, then $\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is jointly convex in $(\bm{\theta},\mathbf{L})\in\Theta\times(L^{\infty})^{n}$ for all $(a,x)\in\mathbb{R}^{2}$ if and only if $f(\bm{\theta},\mathbf{L})$ is jointly convex in $(\bm{\theta},\mathbf{L})$.
@@
    $\widetilde{T}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is convex in $x\in\mathbb{R}$ for all $(\bm{\theta},a,\mathbf{L})\in\Theta\times\mathbb{R}\times(L^{\infty})^{n}$ if and only if the function $x\mapsto 1/(1-\Lambda(x))$ is convex.
@@
The equation (27) holds directly by Theorem 4. Statements (i) and (iii) follow by Theorem 4. Statement (ii) is straightforward. ∎

### v3

@@ 1 Introduction
Beyond the foundational definition and properties of $\Lambda$-ES, which are the topics of Section 3, we proceed to conduct a comprehensive analysis of this new class of risk measures. In Section 4, we obtain a dual representation of $\Lambda$-ES (Theorem 3), offering deeper insights into its theoretical structure and connections to quasi-convex cash-subadditive risk measures. In Section 5, we explore the properties of $\Lambda$-ES in optimization problems, both as an objective function to minimize and as a constraint to impose, and analyze various forms of convexity in relevant reformulations of ES optimization problems. Section 7 concludes the paper.
@@ 2.1 Risk measures
A risk measure is a mapping $\rho:\mathcal{X}\to\overline{\mathbb{R}}$, where $\mathcal{X}$ is the space of random variables where $\rho$ is defined. Below, we list several common properties that a risk measure may satisfy. Their financial interpretation is well documented in the classsic literature (e.g., Artzner et al., 1999; Föllmer and Schied, 2002; Frittelli and Rosazza Gianin, 2002). For some recent developments on these and related properties and their interpretations, see e.g., Castagnoli et al. (2022), Amarante and Liebrich (2024), and Herdegen and Khan (2025).
The risk measure $\rho$ is called a monetary risk measure if it satisfies
@@
  • $L^{1}$-continuity: $\rho(X_{n})\to\rho(X)$ for all $X,X_{1},X_{2}\dots\in\mathcal{X}\cap L^{1}$ and $X_{n}\stackrel{{\scriptstyle L^{1}}}{{\to}}X$ as $n\to\infty$.
In the definition of $L^{1}$-continuity, the space $\mathcal{X}$ may be different from $L^{1}$, such as $L^{\infty}$ or $L^{0}$.
@@ 5.2 Optimization with Lambda ES
We briefly discuss optimization problems with $\Lambda$-ES as a constraint or an objective, which turn out to be closely related to corresponding problems for ES. For $n\in\mathbb{N}$, let $\mathbf{L}=(L_{1},\dots,L_{n})\in(L^{\infty})^{n}$ represent a vector of losses, $\bm{\theta}\in\Theta$ represent a decision variable, where $\Theta$ is a convex set of actions, and $f:\Theta\times\mathbb{R}^{n}\to\mathbb{R}$ represent a loss function. In a typical example in finance, $\mathbf{L}$ represents losses from multiple assets and $\bm{\theta}$ represents a portfolio choice.
First, we are interested in the setting where the decision maker minimizes an objective risk measure $\rho:L^{\infty}\to\overline{\mathbb{R}}$ of the aggregate loss $f(\bm{\theta},\mathbf{L})$, guaranteeing that the $\Lambda$-ES of the total loss does not exceed a pre-specified value $\ell\in\mathbb{R}$. Namely, we consider the following optimization problem:
@@
It is straightforward to show that the constraint in (24) is equivalent an ES constraint under a mild condition.
@@
ES-constrained optimization problems have been studied extensively in finance and operations research (see e.g., Rockafellar and Uryasev, 2002; Krokhmal et al., 2002; Zymler et al., 2013).
A related optimization problem is to minimize $\Lambda$-ES as an objective:
@@
Problems in the form of (25) for ES are also well studied in operations research (see e.g., Zhu and Fukushima, 2009; Cai et al., 2025; Wu et al., 2025). The result below provides general insights into solving (25). Define the mapping $\widetilde{T}^{f}_{\Lambda}:\Theta\times\mathbb{R}\times\mathbb{R}\times(L^{\infty})^{n}\to\overline{\mathbb{R}}$ as
$\widetilde{T}^{f}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})=\mathbb{E}\left[a+\frac{1}{1-\Lambda(x)}(f(\bm{\theta},\mathbf{L})-a)_{+}\right]\vee x.$ (26)
@@
For a right-continuous decreasing function $\Lambda:\mathbb{R}\to[0,1]$ and $\widetilde{T}^{f}_{\Lambda}$ defined in (26),
$\min_{\bm{\theta}\in\Theta}\mathrm{ES}_{\Lambda}(f(\bm{\theta},\mathbf{L}))=\min_{(\bm{\theta},a,x)\in\Theta\times\mathbb{R}\times\mathbb{R}}\widetilde{T}^{f}_{\Lambda}(\bm{\theta},a,x,\mathbf{L}).$ (27)
Moreover,
@@
    $\widetilde{T}^{f}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is convex in $a\in\mathbb{R}$ for all $(\bm{\theta},x,\mathbf{L})\in\Theta\times\mathbb{R}\times(L^{\infty})^{n}$.
@@
    if in addition, $\Lambda$ is not constantly $1$, then $\widetilde{T}^{f}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is jointly convex in $(\bm{\theta},\mathbf{L})\in\Theta\times(L^{\infty})^{n}$ for all $(a,x)\in\mathbb{R}^{2}$ if and only if $f(\bm{\theta},\mathbf{L})$ is jointly convex in $(\bm{\theta},\mathbf{L})$.
@@
    $\widetilde{T}^{f}_{\Lambda}(\bm{\theta},a,x,\mathbf{L})$ is convex in $x\in\mathbb{R}$ for all $(\bm{\theta},a,\mathbf{L})\in\Theta\times\mathbb{R}\times(L^{\infty})^{n}$ if and only if the function $x\mapsto 1/(1-\Lambda(x))$ is convex.
@@
Equation (27) and statements (i) and (iii) follow from Theorem 4. Statement (ii) is straightforward to verify. ∎
Although the results in Propositions 45 are quite simple, they illustrate that $\mathrm{ES}_{\Lambda}$-based optimization problems share many similar features to ES-based ones, which are widely studied.